find the perimeter of a triangle with sides measuring 4 3/4 in 3 5/8 in and 5 11/16 in

thank you

Answer by sarah_adam(201)

perimeter of a triangle = a+b+c
where a, b and c are the sides of the traiangle
here a = 4 3/4 in; b = 3 5/8 in ; c = 5 11/16 in
converting into improper fractions a = 19/4 in ;b = 29/8 in; c = 81/16 in
therfore perimeter = 19/4+29/8+99/16
(19*4 + 29*2 + 81*1)/16
233/16 = 16 7/13 in